# CADSpotting: Robust Panoptic Symbol Spotting on Large-Scale CAD Drawings

### v2

@@ 4 Experiments
4.1 LS-CAD Dataset
We introduce LS-CAD, a new large-scale CAD dataset consisting of 50 floorplans from extensive buildings, such as campuses and office buildings. Each CAD drawing in LS-CAD covers an area of at least 1,000 square meters, with the number of primitives ranging from approximately 2,900 to over ten thousand. Table.1 showcases five large-scale CAD drawings from our dataset, providing detailed information such as area, number of primitives and instance count, which underscore the large-scale nature of LS-CAD.
Additionally, the floorplans in LS-CAD include a wide variety of complex symbols, such as doors, windows, walls, stairs, elevators, different types of furniture, and parking spaces. We provide fine-grained annotations consistent with the standards of the FloorPlanCAD [5] dataset. A portion of LS-CAD will be publicly available under a license waiver to support further research in panoptic symbol spotting for large-scale CAD drawings.
@@ 4.2 Experimental Settings
Implementation Details.
We use the train split of the FloorPlanCAD dataset to train our model. The FloorPlanCAD dataset comprises approximately 15,000 drawings, annotated across 35 symbol categories, including 30 things and 5 stuff classes. For dense point sampling, we set a fixed distance . Model training is performed on a machine with 8 NVIDIA A100 GPUs, using a batch size of 2 per GPU for 512 epochs. We use the AdamW [19] optimizer with a learning rate of and a weight decay of 0.05. Data augmentation techniques include random horizontal flipping with a probability of , global rotation and scaling transformations, translation along the x and y axes and color normalization to . We set loss weight as . Additionally, we set as the number of semantic labels, and both the top- value and are set to 220. We use as the step size of Sliding Window Aggregation on LS-CAD dataset.
Metrics.
Following the approach of [6], we assess the performance of our model using multiple metrics. For semantic symbol spotting, we use F1 score, weighted F1 score (wF1), and mean Intersection over Union (mIoU). For instance symbol spotting, we employ AP50, AP75, and mean Average Precision (mAP). For panoptic symbol spotting, we utilize Panoptic Quality (PQ), Segmentation Quality (SQ), and Recognition Quality (RQ). PQ, in particular, is a comprehensive metric combining SQ and RQ to evaluate model performance in semantic and instance segmentation tasks. Further details on metric formulation are available in [6].
@@ 4.3 Quantitative Evaluation
To demonstrate the effectiveness of our approach, we compare it against SOTA and concurrent methods on the FloorPlanCAD [5] and LS-CAD datasets. A detailed comparison of semantic, instance, and panoptic symbol spotting tasks is provided in the following paragraphs.
Semantic Symbol Spotting.
@@
As shown in Table.2, we compare our method with PanCADNet [5], CADTransformer [6], GAT-CADNet [39], SymPoint [16] and the concurrent SymPoint-V2 [17]. Our CADSpotting approach achieves the highest performance in both the F1 and wF1 metrics, outperforming SymPoint by 13.6% and SymPoint-V2 by 8.6% in mIoU.
Instance Symbol Spotting.
@@
We also evaluate our method against DINO [37], SymPoint and SymPoint-V2 for instance symbol spotting, using the bounding box maximization of predicted mask to compute the AP metric, consistent with SymPoint. As illustrated in Table.3, our method achieves superior results in the AP50, AP75 and mAP metrics, with improvements of 0.9% AP50, 8.4% AP75 and 8.9% mAP over SymPoint-V2.
Panoptic Symbol Spotting.
@@
We then compare our method with the same SOTA and concurrent methods used in the semantic symbol spotting comparison. The results, presented in Table.4, show that our approach surpasses SymPoint and other prior methods across all metrics, including Total/Thing/Stuff PQ, SQ and RQ. Our method achieves results comparable to the concurrent SymPoint-V2, with only a slight performance gap of no more than 1.2% , except for RQ in the Stuff type, where our method outperforms. Additionally, our approach offers superior generalization across diverse CAD primitives, thanks to its dense point sampling based feature representation.
@@
We also compare CADSpotting with SymPoint on our LS-CAD dataset. Since the original SymPoint cannot directly process large-scale CAD drawings, we first apply straightforward block partitioning, using the same size as drawings the FloorPlanCAD dataset without overlapping between blocks, to divide large-scale CAD drawings into smaller blocks for input into SymPoint, referred to as SymPointBP. The results in Table.7 show that our method with block partitioning (CADSpottingBP) achieves the highest performance in SQ. Our method with the SWA technique (CADSpottingSWA) achieves results comparable to SymPointBP and CADSpottingBP, with an absolute improvement of 2.5% in SQ compared to SymPointBP and a slight decrease of 0.8% compared to CADSpottingBP. However, CADSpottingSWA significantly outperforms SymPointBP in PQ and RQ, with an improvement of at least 20%, demonstrating that our CADSpotting method with the SWA technique can accurately recognize instances of CAD symbols in large-scale CAD drawings.
@@
4.4 Qualitative Comparison
We perform qualitative comparisons with SymPoint on the FloorPlanCAD dataset. Fig.4 and Fig.5 show that our method achieves accurate and robust semantic and panoptic symbol spotting, even in challenging scenarios involving walls, complex or overlapping symbols, and uncommon furniture symbols. This demonstrates the effectiveness of our proposed dense point sampling based primitive features. Additionally, Fig.6 presents our semantic and panoptic symbol spotting results on a large-scale CAD drawing of an office building with 28,193 primitives and 1,185 instances from our LS-CAD dataset. These results demonstrate that our CADSpotting enhanced by SWA technique, accurately identifies semantic and instance information in large-scale CAD drawings.
4.5 Ablation Studies
Feature Learning.
We compare our dense point sampling based primitive feature learning method with the approach in SymPoint, which uses handcrafted features to represent each individual primitive. SymPoint employs Point Transformer(PTv1) [38] as its backbone network and incorporates hierarchical multi-resolution primitive features to leverage intermediate features, enhancing the decoder’s performance. To evaluate the impact of different point cloud analysis methods, we first replace SymPoint’s backbone with PTv3 [34], maintaining the same decoder architecture as ours to ensure a fair comparison of feature learning techniques in panoptic symbol spotting. As shown in Table.6, with the same decoder configuration, simply switching to PTv3 results in a 1.8% improvement in PQ. When we apply our proposed primitive feature learning method, performance further increases by 8.7% in PQ, demonstrating the significant performance gains achieved by our dense point sampling based feature learning approach.
| Methods | Backbone | PQ | SQ | RQ |
| Handcrafted Features | PTv1 | 78.4 | 88.5 | 88.6 |
| Handcrafted Features | PTv3 | 80.2 | 90.2 | 89.0 |
| Ours | PTv3 | 88.9 | 95.6 | 93.0 |
| Method | PQ | SQ | RQ |
| SymPoint + Block Partitioning | 33.5 | 84.6 | 39.6 |
| Ours + Block Partitioning | 50.1 | 87.9 | 56.6 |
| Ours + SWA | 54.3 | 87.1 | 61.8 |
Pooling.
We further evaluate the necessity and impact of different pooling methods on our approach, using a baseline configuration without a pooling layer. As shown in Table 5, max pooling improves performance by 11.5%, average pooling by 13.0%, and mixed pooling by 14.3% in PQ. Based on these results, we select mixed pooling for our CADSpotting approach.
Partitioning.
Table.7 shows that our method with the SWA technique (CADSpottingSWA) yields results comparable to our method using block partitioning (CADSpottingBP), with only a slight decrease of 0.8% in SQ. However, CADSpottingSWA surpasses CADSpottingBP in PQ and RQ by a margin, highlighting the effectiveness of the Sliding Window Aggregation technique in panoptic symbol spotting for large-scale CAD drawings.
5 Conclusion
We have introduced CADSpotting, a novel approach for panoptic symbol spotting in large-scale CAD drawings. Our method utilizes dense point sampling within a unified point cloud processing framework to effectively learn robust feature representations of CAD primitives. To address the challenges of precise segmentation in complex, large-scale floorplans, we propose a Sliding Window Aggregation technique, which combines weighted voting with Non-Maximum Suppression to achieve precise symbol detection and segmentation. Additionally, we present LS-CAD, a new large-scale dataset containing 50 annotated CAD floorplans from diverse building types, providing a valuable benchmark for the CAD symbol spotting task. Experimental results on both the LS-CAD and FloorPlanCAD datasets demonstrate the effectiveness, scalability, and robustness of our approach.
Despite these advancements, one limitation of our method is its panoptic symbol spotting performance on large-scale CAD drawings, which currently lags behind its performance on smaller-scale datasets like FloorPlanCAD. Addressing this gap is a focus of our ongoing and future work. We also plan to extend our approach to joint symbol spotting and 3D model generation across multiple building layers, moving towards the goal of fully automated 3D building reconstruction from CAD drawings.
Supplementary Material
Appendix A LS-CAD Dataset
In Figure 8, we present visualizations of several samples from the LS-CAD dataset to highlight the diversity of the proposed dataset. The top panel illustrates Office 3, as described in the main text, which spans a floor area of 2,994 and includes 16,803 primitives and 559 instances. The bottom panel depicts Hotel 1, also described in the main text, which spans a floor area of 1,193 and includes 26,709 primitives and 335 instances.
Appendix B Quantitative Evaluation
Detailed Ablation Study Results
We present a comprehensive evaluation of panoptic quality (PQ), segmentation quality (SQ), and recognition quality (RQ) for each class, comparing different feature learning methods. Our final approach achieves superior performance across the majority of classes, particularly excelling in commonly used categories such as door classes, window classes, and wall classes. The detailed results are shown in Table 9.
Appendix C Qualitative Evaluation
We provide visual comparisons of prediction results on different datasets. For FloorplanCAD, we compare the ground truth (GT) and predicted results using their primitive colors to intuitively evaluate prediction accuracy, as shown in Figure 9 and 10. For LS-CAD, we present the results of semantic and panoptic predictions, highlighting the performance of the method in practical scenarios, as shown in Figure 11 and 12.
Appendix D Inference Time
We conducted a test and comparison of inference speeds for different methods on the FloorplanCAD test set using a single A6000 GPU. The results are presented in Table 8.
| Method | Backbone | Pooling | Speed |
| Handcrafted Features | PTv1 | 83ms | |
| Handcrafted Features | PTv3 | 60ms | |
| Ours | PTv3 | 133ms | |
| Ours | PTv3 | 90ms |
Appendix E Slding Window Implementation Details
In the matrix-NMS [32] process for sliding window aggregation, computing the Intersection over Union (IoU) between all pairs of instance proposal masks is essential for constructing the IoU matrix. However, the number of instance proposals grows substantially with the size of the input drawing. If a dense matrix representation is used, both the time and memory complexities of this operation become prohibitively high.
To mitigate this issue, we observe that only a small fraction of proposal pairs exhibit nonzero overlap, as most proposals belong to different inference windows. Figure 7 visualizes the sparsity structure of the IoU matrix during sliding window aggregation. By leveraging this sparsity, we reimplemented the IoU computation using sparse matrix routines, significantly reducing computational cost and memory usage without sacrificing accuracy.

### v3

@@ 4 Experiments
4.1 The Proposed LS-CAD Dataset
We introduce LS-CAD, a new large-scale CAD dataset consisting of 50 floorplans from extensive buildings, such as campuses and office buildings. Each CAD drawing in LS-CAD covers an area of at least 1,000 square meters, with the number of primitives ranging from approximately 2,900 to over ten thousand. Table.1 showcases 3 large-scale CAD drawings from our dataset, providing detailed information such as area, number of primitives and instance count, which underscore the large-scale nature of LS-CAD.
Additionally, the floorplans in LS-CAD include a wide variety of complex symbols, such as doors, windows, walls, elevators and parking spaces. We provide fine-grained annotations consistent with the standards of the FloorPlanCAD [5] dataset. A portion of LS-CAD will be publicly available under a license waiver to support further research in panoptic symbol spotting for large-scale CAD drawings.
@@ 4.2 Experimental Settings
Implementation Details. The FloorPlanCAD dataset comprises approximately 15,000 drawings, annotated across 35 symbol categories, including 30 things and 5 stuff classes. To ensure fair comparison, we utilize the same training subset from the FloorPlanCAD dataset as used by SymPoint. We train our model on a machine with 8 NVIDIA A100 GPUs, using a batch size of 2 per GPU for 512 epochs. For dense point sampling, we set a fixed distance . We use the AdamW [19] optimizer with a learning rate of and a weight decay of 0.05. Data augmentation techniques include random horizontal flipping with a probability of , global rotation and scaling transformations, translation along the x and y axes and color normalization to . We set loss weight as . Additionally, we set as the number of semantic labels, and both the top- value and are set to 220. We use as the step size of SWA on LS-CAD dataset.
Metrics. Following the approach of [6], we assess the performance of our model using multiple metrics. For semantic symbol spotting, we use F1 score, weighted F1 score (wF1), and mean Intersection over Union (mIoU). For instance symbol spotting, we employ AP50, AP75, and mean Average Precision (mAP). For panoptic symbol spotting, we utilize Panoptic Quality (PQ), Segmentation Quality (SQ), and Recognition Quality (RQ). PQ is a comprehensive metric combining SQ and RQ to evaluate model performance in semantic and instance segmentation tasks. Further details on metric formulation are available in [6].
@@ 4.3 Quantitative Evaluation
To demonstrate the effectiveness of our approach, we compare it against SOTA methods on the FloorPlanCAD [5] and LS-CAD datasets. A detailed comparison of semantic, instance, and panoptic symbol spotting tasks is provided in the following paragraphs.
@@
Semantic Symbol Spotting. As shown in Table.2, we compare our method with PanCADNet [5], CADTransformer [6], GAT-CADNet [40], and SymPoint [16]. Our CADSpotting approach achieves the highest performance in both the F1 and wF1 metrics, outperforming SymPoint by 13.6% in mIoU.
@@
Instance Symbol Spotting. We also evaluate our method against DINO [38] and SymPoint for instance symbol spotting, using the bounding box maximization of predicted mask to compute the AP metric, consistent with SymPoint. As shown in Table.3, our method achieves superior results in the AP50, AP75 and mAP metrics, with improvements of 5.9% AP50, 8.4% AP75 and 8.9% mAP over SymPoint.
@@
Panoptic Symbol Spotting. We then compare our method with the same methods used in the semantic symbol spotting comparison. The results, presented in Table.4, show that our approach surpasses SymPoint and other prior methods across all metrics, including Total/Thing/Stuff PQ, SQ and RQ. Additionally, our approach offers superior generalization across diverse CAD primitives, thanks to its dense point sampling based feature representation.
@@
We also compare the effectiveness of Sliding Window Aggregation (SWA) and Block Partitioning (BP) methods when applied to the CADSpotting model on the LS-CAD dataset. We apply BP method using the same size as drawings the FloorPlanCAD dataset without overlapping between blocks, to divide large-scale CAD drawings into smaller blocks. When employing the BP method, we ensure that each primitive is uniquely represented by retaining only those whose starting points fall within the block. The results in Table 5 indicate that, compared to the BP method, SWA achieves the highest performance on PQ, with an absolute improvement of 10.4%, although there is a slight decrease of 1.0% on SQ. The result shows that our CADSpotting method with the SWA technique can accurately recognize instances of CAD symbols in large-scale CAD drawings.
4.4 Qualitative Comparison
We perform qualitative comparisons with SymPoint on the FloorPlanCAD dataset. Fig.5 shows that our method achieves accurate and robust panoptic symbol spotting, even in challenging scenarios involving walls, complex or overlapping symbols, and uncommon furniture symbols. This demonstrates the effectiveness of our proposed dense point sampling based primitive features. Additionally, Fig.6 presents our semantic and panoptic symbol spotting results on a large-scale CAD drawing of an office building with 28,193 primitives and 1,185 instances from our LS-CAD dataset. These results demonstrate that our CADSpotting enhanced by SWA technique, accurately identifies semantic and instance information in large-scale CAD drawings.
| Method | PQ | SQ | RQ |
| Ours + BP | 42.8 | 94.8 | 45.2 |
| Ours + SWA | 53.2 | 93.8 | 56.6 |
To evaluate model generalizability and demonstrate the utility of LS-CAD dataset, we also train models on a combined dataset consisting of FloorPlanCAD and LS-CAD. Comprehensive experimental results and analyses are provided in the supplementary material.
| Methods | Backbone | PQ | SQ | RQ |
| Handcrafted Features | PTv1 | 78.4 | 88.5 | 88.6 |
| Handcrafted Features | PTv3 | 80.2 | 90.2 | 89.0 |
| Ours | PTv3 | 88.9 | 95.6 | 93.0 |
| Method | PQ | SQ | RQ |
| w/o Color Feature | 85.8 | 92.5 | 92.7 |
| Color Feature | 88.9 | 95.6 | 93.0 |
4.5 Ablation Studies
In this section, we present ablation studies examining the effects of feature learning and RGB information. Additional ablation studies, which explore various pooling strategies and evaluate performance across all classes, are provided in the supplementary material.
Feature Learning. We compare our dense point sampling based primitive feature learning method with the approach in SymPoint, which uses handcrafted features to represent each individual primitive. SymPoint employs Point Transformer(PTv1) [39] as its backbone network and incorporates hierarchical multi-resolution primitive features to leverage intermediate features, enhancing the decoder’s performance. To evaluate the impact of different point cloud analysis methods, we first replace SymPoint’s backbone with PTv3 [35], maintaining the same decoder architecture as ours to ensure a fair comparison of feature learning techniques in panoptic symbol spotting. As shown in Table.6, with the same decoder configuration, simply switching to PTv3 results in a 1.8% improvement in PQ. When we apply our proposed primitive feature learning method, performance further increases by 8.7% in PQ, demonstrating the significant performance gains achieved by our dense point sampling based feature learning approach.
RGB Color Information. We compare the performance of CADSpotting through the inclusion versus exclusion of color information in the FloorPlanCAD dataset. Table.7 shows that adding RGB information improves the PQ metric by 3.1%.
5 Conclusion
We have introduced CADSpotting, a novel method for panoptic symbol spotting in large-scale CAD drawings. Our approach uses dense point sampling within a unified point cloud framework to represent CAD primitives, combined with a Sliding Window Aggregation technique that integrates weighted voting and Non-Maximum Suppression for precise segmentation. We also present LS-CAD, a dataset of 50 annotated floorplans from diverse building types. Experiments on LS-CAD and FloorPlanCAD datasets have confirmed the effectiveness and scalability of our method, with successful applications in automated 3D reconstruction demonstrating its practical value.
Our approach currently faces limitations due to the dataset’s primary emphasis on residential and commercial buildings, which restricts its applicability to industrial and cultural contexts. Additionally,it may struggle to capture significant stylistic variations among CAD symbols. Addressing these challenges through the integration of in-context learning may help enhance annotation efficiency and extend applicability. This is left as our future work.
Supplementary Material
| Method | Training Dataset | Test on FloorPlanCAD | Test on Cross-dataset | ||||
| PQ | SQ | RQ | PQ | SQ | RQ | ||
| SymPoint [16] | FloorPlanCAD | 83.3 | 91.4 | 91.1 | 33.2 | 83.0 | 39.9 |
| SymPoint | Cross-dataset | 79.0 | 89.2 | 88.6 | 57.0 | 84.0 | 67.8 |
| CADSpotting | FloorPlanCAD | 88.9 | 95.6 | 93.0 | 42.8 | 94.8 | 45.2 |
| CADSpotting | Cross-dataset | 87.3 | 95.4 | 91.4 | 69.9 | 91.6 | 76.3 |
| CADSpotting(SWA) | FloorPlanCAD | / | 53.2 | 93.8 | 56.6 | ||
| CADSpotting(SWA) | Cross-dataset | / | 75.5 | 84.9 | 88.9 | ||
6 More Details on the LS-CAD Dataset
In Fig.8, we present visualizations of several samples from the LS-CAD dataset to highlight the diversity of the proposed dataset. The top panel illustrates Office 3, which spans a floor area of 2,994, including 16,803 primitives and 559 instances. The bottom panel depicts Hotel 1, which spans a floor area of 1,193, consisting of 26,709 primitives and 335 instances.
To demonstrate the generalization capability of our model and highlight the value of the LS-CAD dataset, we implement a cross-dataset joint training paradigm. We randomly partition complete large-scale CAD drawings into three subsets: 80% for training, 10% for validation, and 10% for testing. The division of the LS-CAD dataset strictly adheres to the standardized data splitting protocol established in FloorPlanCAD. The training and validation sets from both datasets are subsequently integrated to construct a cross-dataset collaborative training environment. As shown in Table.8, the cross-dataset joint training strategy significantly improves model performanc on the LS-CAD test set, although with slight performance degradation on the original FloorPlanCAD test set. This demonstrates the necessity of the proposed LS-CAD dataset for provding more diverse CAD representations. Notably, CADSpotting+SWA method reaches the best performance of 75.5 on the PQ metric. After cross-dataset joint training, the PQ score for CADSpotting decreases by only 1.2% on FloorPlanCAD, whereas it drops by 4.3% on SymPoint. This indicates that our model exhibits stronger generalization capabilities.
7 Quantitative Evaluation
Ablation Study on Feature Learning.
We present a comprehensive evaluation of panoptic quality (PQ), segmentation quality (SQ), and recognition quality (RQ) for each class, comparing different feature learning methods. Our final approach achieves superior performance across the majority of classes, particularly excelling in commonly used categories such as door classes, window classes, and wall classes. The detailed results are shown in Table.10.
Ablation Study on Pooling.
We further evaluate the necessity and impact of different pooling methods on our approach, using a baseline configuration without a pooling layer. As shown in Table.9, max pooling improves performance by 11.5%, average pooling by 13.0%, and mixed pooling by 14.3% in PQ. Based on these results, we select mixed pooling for our CADSpotting approach.
8 Qualitative Evaluation
We provide comparisons of prediction results on different datasets. For FloorPlanCAD, we compare the ground truth and prediction using their primitive colors to intuitively evaluate prediction accuracy, as shown in Fig.9, Fig.10 and Fig.13. For LS-CAD, we present the results of semantic and panoptic predictions, highlighting the performance of the method in practical scenarios, as shown in Fig.11 and Fig.12.
9 Sliding Window Implementation Details
The Algorithm.1 details the implementation process of our SWA. In the matrix-NMS [33] process for sliding window aggregation, computing the Intersection over Union (IoU) between all pairs of instance proposal masks is essential for constructing the IoU matrix. However, the number of instance proposals grows substantially with the size of the input drawing. If a dense matrix representation is used, both the time and memory complexities of this operation become prohibitively high.
To mitigate this issue, we observe that only a small fraction of proposal pairs exhibit nonzero overlap, as most proposals belong to different inference windows. Fig. 7 visualizes the sparsity structure of the IoU matrix during sliding window aggregation. By leveraging this sparsity, we reimplemented the IoU computation using sparse matrix routines, significantly reducing computational cost and memory usage without sacrificing accuracy.
10 Automated 3D Interior Reconstruction
Based on the spotting results generated by CADSpotting, we present more 3D reconstruction renderings in Fig.14. By systematically optimizing the parameters of predefined instance components using instance segmentation data and primitive spatial coordinates, we achieve accurate 3D model reconstruction through component-level geometric adaptation
For wall polygon extraction, we employ an innovative SVG-PNG conversion strategy to differentiate floor areas from walls: 1) Merge adjacent endpoints and remove duplicate segments to simplify SVG paths. 2) Convert vectors to 8K resolution binarized PNG while maintaining coordinate alignment. 3) Identify the largest connected component as the floor through connected component analysis, with secondary components as candidate walls. 4) Re-vectorize boundaries through coordinate mapping to generate closed wall polygons.
The proposed method exhibits notable robustness against recognition inaccuracies – incomplete walls are automatically reconciled during stage 1, while small noise particles are filtered in stage 3 using area thresholds.
For doors and windows, we employ category-specific parameterization strategies. Doors are systematically categorized into four distinct subtypes (single/double/sliding/folding) based on semantic labels, with positions, orientations, and pivot points determined through linear-arc spatial relationships.
Windows are grouped via union-find algorithms, calculating average lines of co-grouped segments as representative lines, then selecting the segment closest to the representative line within each group as the window instance position.
The final reconstruction leverages Blender’s geometry nodes for procedural modeling: 1) Extrude wall polygons to 3D volumes using floorplan height attributes. 2) Instance-place door/window assemblies at infered positions. 3) Apply floor based on origin CAD drawing.
This integrated approach achieves automated 3D reconstruction with exceptional efficiency, transforming raw CAD data into structurally complete 3D interior models within minutes through streamlined procedural generation.